heres another.
i dont get it when there on each side of the equal sign
-2(x+11)=-x+10
oh and
3(y+6)=2y+14
and i have to check which i also dont understand!

Answer by checkley77(12844)

-2(x+11)=-x+10
-2x-22=-x+10
-2x+x=10+22
-x=32 answer.
x=-32
Proof:
-2(-32+11)=-(-32)+10
-2*-21=32+10
42=42
--------------------------------
3(y+6)=2y+14
3y+18=2y+14
3y-2y=14-18
y=-4 answer.
Proof:
3(-4+6)=2(-4)+14
3*2=-8+14
6=6